How do I solve the square root of 6x-3 is equal to square root of 8w-15? I have to solve for X
Answer by rothauserc(4718)

square root(6x-3) = square root(8w-15)
square both sides of the =
6x-3 = 8w-15
6x = 8w-12
x = (8/6)w -2
x = (4/3)w -2